Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

night-report-2015-10-11


Observers: Joe Lyman, Heather Campbell
TAT: Janet Chen, Annalisa De Cia
Data reducers: Morgan Fraser

Conditions good at start of night, only thin cloud on the horizon. Wind <10m/s humidity ~50%
seeing <1'' at the start of the night, ended with cloud

23.05 Started Sky flats gri

23.35 Standard l175

23.55 TCP17 classification in G11+G16 - G16 internal flat not worked, still on sky, the TO is reporting the problem. - looks very red, lots of absorption features and Halpha in emission

23.10 AS15qy - Type Ia ~5 days before peak - not in marshall yet, added from ATEL as good constraint on non detection and bright

23.25 PS15cgd - repeated from last night, observed for 1200s in good conditions - spectrum looks very similar. - can also get a spectrum of the other galaxy. 

24.50 Switch to SOFI

01.00  image analysis

01.15 Standard Fiege100 seeing ~<1''

01.35 LSQ15adm JHK imaging seeing ~0.9'' in J,H band and seeing ~1.1'' in k band

02.50 AS15qc Telluric 

03.00 AS15qc BG+RG 

05.30 AS15qc Telluric 

05.35 AS15qc imaging seeing ~1.4''

05.50 AS14Ko imaging seeing ~1.4''

06.05 changed to EFOSC

06.25 Standard

06.45 ASASN-14ha R band 18*100s imaging seeing <1.1'' thin clouds on horizon

07.35 PSNJ0205 45mins 1000s G11+G16 some thin cloud coming in during the G16 exposure

Cloud much thicker couldn't see any stars in acquisition images and clouds in red. Mirror closed. Waiting to see if improves.

Conditions stayed bad and had to close telescope

Comments